You invest $280 in a mutual fund today that pays 9.10 percent interest annually. How long will it take to double your money? (If you solve this problem with algebra round intermediate calculations to 6 decimal places, in all cases round your final answer to 0 decimal places.)
Answer #1

future amount = amount invested*(1+r)^n

here,

future amount = 280*2 =>560.

amount invested = 280

r = 9.10%=>0.091.

n = to be found out.

560 = 280 *(1.091)^n

=>2 = 1.09^n

apply logarithms

=>log 2 = log (1.09)^n

=>log 2 = n log (1.09)

=>log 2 / log 1.09 = n

=>8.04323173

=>8 years.

It will take 8 years to double the money.